A tire manufacturer would like to estimate the average tire life of its new​ all-season light truck tire in terms of how many miles it lasts. Determine the sample size needed to construct a 95​% confidence interval with a margin of error equal to 1,500 miles. Assume the standard deviation for the tire life of this particular brand is 10,000 miles.

Answer: The sample size needed is 171.

Above are the questions and answers. Please provide all Excel formulas (not calculations) that get to these answers.
Answer #1

Margin of Error = E= 1500

value of z at 95% confidence interval = 1.96

Substituting the values in the formula of Margin of error

  

In excel

=1.96 * 10000

In excel

=19600 / 1500

In excel

=13.067 * 13.067

ANS: Sample size needed - 171